# Global solutions to cross-diffusion systems with independent advections in one dimension

### 20 Mar 2026 v1

@@ 3.4. A conservation law applicable for $\alpha\in(0,2]$
$\begin{split}\frac{2\,u_{\varepsilon}\,s_{\varepsilon}}{s_{\varepsilon}^{2}}\,&\left[\partial_{x}u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})\right]-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\left[\partial_{x}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+s_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})\right]=\\ &=\frac{s_{\varepsilon}}{s_{\varepsilon}^{2}}\,\partial_{x}u_{\varepsilon}^{2}\,\partial_{x}p(s_{\varepsilon})+\frac{s_{\varepsilon}}{s_{\varepsilon}^{2}}\,u_{\varepsilon}^{2}\,\partial_{x}^{2}p(s_{\varepsilon})-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\\ &=\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\right)\,\partial_{x}p(s_{\varepsilon})+\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}^{2}p(s_{\varepsilon})=\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}p(s_{\varepsilon})\right)=\frac{1}{\alpha}\,\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}\right),\end{split}$ (3.21)
@@
$-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}(u_{\varepsilon}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]))=$
@@
$\begin{split}2\,\varepsilon\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial^{2}_{x}u_{\varepsilon}-&\varepsilon\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial^{2}_{x}s_{\varepsilon}=\\ &=2\,\varepsilon\,\partial_{x}\left(\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}u_{\varepsilon}\right)-\varepsilon\,\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}\right)-2\,\varepsilon\,s_{\varepsilon}\,\left(\frac{\partial_{x}u_{\varepsilon}\,s_{\varepsilon}-u_{\varepsilon}\,\partial_{x}s_{\varepsilon}}{s_{\varepsilon}^{2}}\right)^{2}.\end{split}$ (3.24)
@@
$\begin{split}\left[\int_{-\infty}^{x}\partial_{t}s_{\varepsilon}\mathop{}\!\mathrm{d}y\right]\,\partial_{x}s_{\varepsilon}^{\alpha}=\frac{1}{\alpha}|\partial_{x}s_{\varepsilon}^{\alpha}|^{2}+u_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}\,&\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\,+\\ &+s_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]+\varepsilon\,\partial_{x}s_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha},\end{split}$ (3.28)
@@
$\left[\int_{-\infty}^{x}\partial_{t}s_{\varepsilon}\mathop{}\!\mathrm{d}y\right]\,\partial_{x}s_{\varepsilon}^{\alpha}\to\left[\int_{-\infty}^{x}\partial_{t}s\mathop{}\!\mathrm{d}y\right]\,\partial_{x}s^{\alpha}\mbox{ in }\mathcal{D}((0,T)\times\mathbb{R}).$ (3.30)
@@ 4.1. Young measures
For each $(t,x)$ we define the measure $\pi_{t,x}$ on $\mathbb{R}\times\mathbb{R}$ to be the Young measure (see [51] for the original paper by Young, who addressed the lack of strong convergence in problems of the calculus of variations; [5] for the work of Ball, who introduced this concept into the study of PDEs; and [39, Chapter 6] and [44, Chapter 4] for more comprehensive treatments) of the sequence
@@
To say that $\pi_{t,x}$ is the Young measure of sequence in (4.1) means that:
@@
According to [39, Theorem 6.2], the Young measure of a sequence $\{z_{\varepsilon}\}$ exists whenever $\sup_{\varepsilon\in(0,1)}\int_{0}^{T}\int_{\mathbb{R}}g(|z_{\varepsilon}|)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t<\infty$ for some $g:[0,\infty)\to[0,\infty]$ continuous, nondecreasing and $\lim_{t\to\infty}g(t)=\infty$ which is clearly satisfied in our case by (E3) in Proposition 3.1 and $\frac{u_{\varepsilon}}{{s_{\varepsilon}}}\leq 1$.
We will also need some information when the sequence $\Big\{f\Big(\frac{u_{\varepsilon}}{{s_{\varepsilon}}}-\frac{u}{{s}},\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}\Big)\Big\}_{\varepsilon\in(0,1)}$ is only bounded in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$. This is the argument taken from [27, Lemma 2.1] simplified to our setting.
@@
$\widetilde{f}\geq\int_{\mathbb{R}^{2}}f(\lambda_{1},\lambda_{2})\mathop{}\!\mathrm{d}\pi_{t,x}(\lambda_{1},\lambda_{2})\mbox{ in }\mathcal{D}((0,T)\times\mathbb{R}).$ (4.3)
@@ 4.2. Div-Curl Lemma and compensated compactness
The idea is now to find some constraints on $\pi_{t,x}$. The main tool will be a version of Div-Curl Lemma [25, p.54] (see also [38, 47] for the original references) whose original statement is recalled below:
@@
Let $\{\bf a_{\varepsilon}\}_{\varepsilon\in(0,1)}$, $\{\bf b_{\varepsilon}\}_{\varepsilon\in(0,1)}$ be two sequences converging weakly in $L^{2}_{\operatorname{loc}}(\Omega;\mathbb{R}^{k})$ to some $\bf a$}, $\bf b$} for some $\Omega\subset\mathbb{R}^{k}$. Suppose moreover that $\{\operatorname{div}{\bf a_{\varepsilon}}\}_{\varepsilon\in(0,1)}$ and $\{\operatorname{curl}{\bf b_{\varepsilon}}\}_{\varepsilon\in(0,1)}$ lie in a compact subset of $H^{-1}_{\operatorname{loc}}(\Omega)$. Then, $a_{\varepsilon}\cdot b_{\varepsilon}\to a\cdot b$ as $\varepsilon\to 0$ in the sense of distributions.
@@
with $\{U_{\varepsilon}\}_{\varepsilon\in(0,1)}$, $\{V_{\varepsilon}\}_{\varepsilon\in(0,1)}$, $\{F_{\varepsilon}\}_{\varepsilon\in(0,1)}$, $\{G_{\varepsilon}\}_{\varepsilon\in(0,1)}$ bounded in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ and $\{S_{\varepsilon}\}_{\varepsilon\in(0,1)}$, $\{R_{\varepsilon}\}_{\varepsilon\in(0,1)}$ strongly compact in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, one applies Theorem 4.2 with
@@
We will apply these ideas twice and for the second application we need to slightly relax the assumptions of Theorem 4.2 to allow for situations where the divergence includes a term which is bounded only in $L^{1}((0,T)\times\mathbb{R})$ (note carefully that in two dimensions this space does not embed compactly in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ (but only into $W^{-1,p}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ with $p<2$, see [25, p.7]). Similar term arise in the inviscid limit of hyperbolic conservation laws formulated in terms of the entropy/entropy-flux pairs (see [25, p.59])
$\partial_{t}\Phi(u_{\varepsilon})+\partial_{x}\Psi(u_{\varepsilon})=\varepsilon\,\partial_{x}^{2}\Phi(u_{\varepsilon})-\varepsilon\,\Phi^{\prime\prime}(u_{\varepsilon})\,|\partial_{x}u_{\varepsilon}|^{2},$ (4.4)
where $\Phi$, $\Psi$ are bounded nonlinearities and $\{\varepsilon\,\Phi^{\prime\prime}(u_{\varepsilon})\,|\partial_{x}u_{\varepsilon}|^{2}\}_{\varepsilon\in(0,1)}$ is known to be bounded only in $L^{1}((0,T)\times\mathbb{R})$. In this case, one exploits the fact that the (LHS) of (4.4) is bounded in $W^{-1,p}((0,T)\times\mathbb{R})$ with arbitrarily large $p$ so one gets compactness in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ by interpolation and the embedding of $\mathcal{M}_{\operatorname{loc}}([0,T]\times\mathbb{R})$ into $W^{-1,p}((0,T)\times\mathbb{R})$ with $p<2$. In our case, these ideas are inapplicable to the continuity equations (3.17) and (3.19) because $(E_{\varepsilon}+s_{\varepsilon})\,\partial_{x}p(s_{\varepsilon})$ and $\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}$ are bounded only in $L^{2}((0,T)\times\mathbb{R})$. The presence of derivatives of $s_{\varepsilon}$ does not allow to improve the integrability easily.
Here, we will prove the following result which is clearly not stated in the most general form and interested readers may try adapting it to the problems of their interests. The main idea is that we work with two equations: one involving terms with better regularity and another involving terms with lower regularity (more precisely, the (RHS) in the better equation will be compact in $L^{2}_{\operatorname{loc}}(0,T;H^{-1}_{\operatorname{loc}}(\mathbb{R}))$ rather than in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ only). In proofs based on compensated compactness, we never multiply terms coming from the same equation, so the better regularity can always compensate for the worse regularity. We also remark that since we only want to identify the limits, all the necessary estimates are required to hold only locally in time and space.

### 16 Apr 2026 v2

@@ 3.4. A conservation law applicable for $\alpha\in(0,2]$
$\begin{split}\frac{2\,u_{\varepsilon}\,s_{\varepsilon}}{s_{\varepsilon}^{2}}\,&\left[\partial_{x}u_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+u_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})\right]-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\left[\partial_{x}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})+s_{\varepsilon}\,\partial_{x}^{2}p(s_{\varepsilon})\right]\\ &=\frac{s_{\varepsilon}}{s_{\varepsilon}^{2}}\,\partial_{x}u_{\varepsilon}^{2}\,\partial_{x}p(s_{\varepsilon})+\frac{s_{\varepsilon}}{s_{\varepsilon}^{2}}\,u_{\varepsilon}^{2}\,\partial_{x}^{2}p(s_{\varepsilon})-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}\,\partial_{x}p(s_{\varepsilon})\\ &=\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\right)\,\partial_{x}p(s_{\varepsilon})+\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}^{2}p(s_{\varepsilon})=\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}}\,\partial_{x}p(s_{\varepsilon})\right)=\frac{1}{\alpha}\,\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}^{\alpha}\right),\end{split}$ (3.21)
@@
$-\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}(u_{\varepsilon}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]))$
@@
$\begin{split}2\,\varepsilon\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial^{2}_{x}u_{\varepsilon}-&\varepsilon\,\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial^{2}_{x}s_{\varepsilon}\\ &=2\,\varepsilon\,\partial_{x}\left(\frac{u_{\varepsilon}}{s_{\varepsilon}}\,\partial_{x}u_{\varepsilon}\right)-\varepsilon\,\partial_{x}\left(\frac{u_{\varepsilon}^{2}}{s_{\varepsilon}^{2}}\,\partial_{x}s_{\varepsilon}\right)-2\,\varepsilon\,s_{\varepsilon}\,\left(\frac{\partial_{x}u_{\varepsilon}\,s_{\varepsilon}-u_{\varepsilon}\,\partial_{x}s_{\varepsilon}}{s_{\varepsilon}^{2}}\right)^{2}.\end{split}$ (3.24)
@@
Remark 3.8.
We remark that the restriction $\alpha\leq 2$ is only needed to ensure that $\{g_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$. This is achieved by proving that $\{\partial_{x}s_{\varepsilon}\}_{\varepsilon\in(0,1)}$ is bounded in $L^{1}(0,T;L^{1}_{\operatorname{loc}}(\mathbb{R}))$. If such a bound could be established for $\alpha>2$, the conclusion of Proposition 3.7 would also hold in this case. Furthermore, the proof of Proposition 5.2 presented in Section 5.4 would also be valid, yielding an alternative (and simpler) proof of Theorem 1.3. It is not clear whether the range $\alpha\leq 2$ is optimal, and this question appears to be delicate even for the pure porous medium equation $\partial_{t}s=\partial_{x}^{2}s^{\alpha}$ on $\mathbb{R}$ without advection (see, e.g., [55, 32]). A naive approach to prove that $\partial_{x}s\in L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ in this case is to use the identity
${\alpha\,\kappa}\int_{0}^{t}\int_{\mathbb{R}}s^{\alpha+\kappa-2}\,|\partial_{x}s|^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau=\frac{1}{1+\kappa}\int_{\mathbb{R}}(s(t,x)^{1+\kappa}-s(0,x)^{1+\kappa})\mathop{}\!\mathrm{d}x,$ (3.25)
where $\kappa\in(-1,0]$. By splitting the spatial integral into the regions $|x|\leq 1$ and |x|>1|x|~>~1, the terms on the right-hand side can be controlled using the mass $\int_{\mathbb{R}}s(t,x)\mathop{}\!\mathrm{d}x$ and a sufficiently high moment $\int_{\mathbb{R}}s(t,x)\,|x|^{\sigma}\mathop{}\!\mathrm{d}x$, since $1+\kappa\in(0,1]$. Here we use that, for the porous medium equation, all moments are controlled provided the initial data is bounded and has finite moments of all orders. We conclude that $\partial_{x}s\in L^{2}((0,T)\times\mathbb{R})$ provided $\alpha+\kappa-2<0$, which can be achieved when $\alpha<3$. This does not seem to be optimal, since for the Barenblatt’s solution $B(t,x)=t^{-1/(\alpha+1)}\,F(x\,t^{-1/(\alpha+1)})$ with $F(y)=(C-y^{2})_{+}^{1/(\alpha-1)}$, a direct computation shows that $B(t,x)\in L^{1}((\tau,T)\times\mathbb{R})$ for all $\tau>0$. Regarding our problem (1.4), the argument exploiting (3.25) cannot be adapted, since due to the advection we obtain an additional term on the (RHS) of (3.25), which is roughly $\int_{0}^{t}\int_{\mathbb{R}}s_{\varepsilon}^{\kappa}\,|\partial_{x}s_{\varepsilon}|\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau$. By writing $s_{\varepsilon}^{\kappa}\,|\partial_{x}s_{\varepsilon}|=s_{\varepsilon}^{\frac{\kappa+\alpha-2}{2}}\,|\partial_{x}s_{\varepsilon}|\,s_{\varepsilon}^{\frac{\kappa-\alpha+2}{2}}$, we see that this term can be controlled by $\int_{0}^{t}\int_{\mathbb{R}}s_{\varepsilon}^{\alpha+\kappa-2}\,|\partial_{x}s_{\varepsilon}|^{2}\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}\tau$ only if $\kappa-\alpha+2\geq 0$, i.e. $\alpha\leq 2$.
Remark 3.9.
The same computation as in the proof of Proposition 3.7 shows that for any $\theta>0$ we have
$\partial_{t}\left(\frac{u_{\varepsilon}^{\theta+1}}{s_{\varepsilon}^{\theta}}\right)=\partial_{x}\left(\frac{1}{\alpha}\,\frac{u_{\varepsilon}^{\theta+1}}{s_{\varepsilon}^{\theta+1}}\,\partial_{x}s_{\varepsilon}^{\alpha}+f_{\varepsilon,\theta}\right)+g_{\varepsilon,\theta}+h_{\varepsilon,\theta},$
where $f_{\varepsilon,\theta}$, $g_{\varepsilon,\theta}$, $h_{\varepsilon,\theta}$ are defined as follows:
$f_{\varepsilon,\theta}=\,$ $\frac{u_{\varepsilon}^{\theta+1}}{s_{\varepsilon}^{\theta}}\,\partial_{x}\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\frac{\theta}{\theta+2}\,\frac{u_{\varepsilon}^{\theta+2}}{s_{\varepsilon}^{\theta+1}}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]),$
$g_{\varepsilon,\theta}=\,$ $\theta\,\frac{u_{\varepsilon}^{\theta+1}}{s_{\varepsilon}^{\theta+1}}\,\left(1-\frac{(\theta+1)}{(\theta+2)}\,\frac{u_{\varepsilon}}{s_{\varepsilon}}\right)\,\partial_{x}\big(s_{\varepsilon}\,\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\big)$
$-\theta\,(\theta+1)\,\varepsilon\,s_{\varepsilon}\,\frac{u_{\varepsilon}^{\theta-1}}{s_{\varepsilon}^{\theta-1}}\,\left(\frac{\partial_{x}u_{\varepsilon}\,s_{\varepsilon}-u_{\varepsilon}\,\partial_{x}s_{\varepsilon}}{s_{\varepsilon}^{2}}\right)^{2},$
$h_{\varepsilon,\theta}=\,$ $(\theta+1)\,\varepsilon\,\partial_{x}\left(\frac{u_{\varepsilon}^{\theta}}{s_{\varepsilon}^{\theta}}\,\partial_{x}u_{\varepsilon}\right)-\theta\,\varepsilon\,\partial_{x}\left(\frac{u_{\varepsilon}^{\theta+1}}{s_{\varepsilon}^{\theta+1}}\,\partial_{x}s_{\varepsilon}\right).$
These quantities satisfy the same bounds as in Proposition 3.7. This yields a whole family of conservation laws, which is applicable only for $\alpha\leq 2$, for the same reason as explained in Remark 3.8.
@@
$\begin{split}\left[\int_{-\infty}^{x}\partial_{t}s_{\varepsilon}\mathop{}\!\mathrm{d}y\right]\,\partial_{x}s_{\varepsilon}^{\alpha}=\frac{1}{\alpha}|\partial_{x}s_{\varepsilon}^{\alpha}|^{2}+u_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}\,&\partial_{x}(\mathcal{V}^{1}[u_{\varepsilon},v_{\varepsilon}]-\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}])\\ &+s_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha}\,\partial_{x}\mathcal{V}^{2}[u_{\varepsilon},v_{\varepsilon}]+\varepsilon\,\partial_{x}s_{\varepsilon}\,\partial_{x}s_{\varepsilon}^{\alpha},\end{split}$ (3.29)
@@
$\left[\int_{-\infty}^{x}\partial_{t}s_{\varepsilon}\mathop{}\!\mathrm{d}y\right]\,\partial_{x}s_{\varepsilon}^{\alpha}\to\left[\int_{-\infty}^{x}\partial_{t}s\mathop{}\!\mathrm{d}y\right]\,\partial_{x}s^{\alpha}\mbox{ in }\mathcal{D}^{\prime}((0,T)\times\mathbb{R}).$ (3.31)
@@ 4.1. Young measures
For each $(t,x)$ we define the measure $\pi_{t,x}$ on $\mathbb{R}\times\mathbb{R}$ to be the Young measure (see [60] for the original paper by Young, who addressed the lack of strong convergence in problems of the calculus of variations; [5] for the work of Ball, who introduced this concept into the study of PDEs; and [46, Chapter 6] and [51, Chapter 4] for more comprehensive treatments) of the sequence
@@
To say that $\pi_{t,x}$ is the Young measure of the sequence in (4.1) means that:
@@
According to [46, Theorem 6.2], the Young measure of a sequence $\{z_{\varepsilon}\}$ exists whenever $\sup_{\varepsilon\in(0,1)}\int_{0}^{T}\int_{\mathbb{R}}g(|z_{\varepsilon}|)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}t<\infty$ for some $g:[0,\infty)\to[0,\infty]$ continuous, nondecreasing and $\lim_{t\to\infty}g(t)=\infty$ which is clearly satisfied in our case by (E3) in Proposition 3.1 and $\frac{u_{\varepsilon}}{{s_{\varepsilon}}}\leq 1$.
We will also need some information when the sequence $\Big\{f\Big(\frac{u_{\varepsilon}}{{s_{\varepsilon}}}-\frac{u}{{s}},\partial_{x}s_{\varepsilon}^{\alpha}-\partial_{x}s^{\alpha}\Big)\Big\}_{\varepsilon\in(0,1)}$ is only bounded in $L^{1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$. This is the argument taken from [31, Lemma 2.1] simplified to our setting.
@@
$\widetilde{f}\geq\int_{\mathbb{R}^{2}}f(\lambda_{1},\lambda_{2})\mathop{}\!\mathrm{d}\pi_{t,x}(\lambda_{1},\lambda_{2})\mbox{ in }\mathcal{D}^{\prime}((0,T)\times\mathbb{R}).$ (4.3)
@@ 4.2. Div-Curl Lemma and compensated compactness
The idea is now to find some constraints on $\pi_{t,x}$. The main tool will be a version of Div-Curl Lemma [29, p.54] (see also [45, 56] for the original references) whose original statement is recalled below:
@@
Let $\{\bf a_{\varepsilon}\}$, $\{\bf b_{\varepsilon}\}$ be two sequences converging weakly in $L^{2}_{\operatorname{loc}}(\Omega;\mathbb{R}^{k})$ to some $\bf a$}, $\bf b$} for some $\Omega\subset\mathbb{R}^{k}$. Suppose moreover that $\{\operatorname{div}{\bf a_{\varepsilon}}\}$ and $\{\operatorname{curl}{\bf b_{\varepsilon}}\}$ lie in a compact subset of $H^{-1}_{\operatorname{loc}}(\Omega)$. Then, $a_{\varepsilon}\cdot b_{\varepsilon}\to a\cdot b$ as $\varepsilon\to 0$ in the sense of distributions.
@@
with $\{U_{\varepsilon}\}$, $\{V_{\varepsilon}\}$, $\{F_{\varepsilon}\}$, $\{G_{\varepsilon}\}$ bounded in $L^{2}_{\operatorname{loc}}((0,T)\times\mathbb{R})$ and $\{S_{\varepsilon}\}$, $\{R_{\varepsilon}\}$ strongly compact in $H^{-1}_{\operatorname{loc}}((0,T)\times\mathbb{R})$, one applies Theorem 4.2 with
@@
We will apply these ideas twice and for the second application we need to slightly relax the assumptions of Theorem 4.2. The following variant is proved in [21].